# Quantum correction to the orbital Hall effect

### v1

@@ Abstract
How is one to interpret the presence of band-diagonal position matrix elements, given that Bloch electrons are delocalized and their position is undefined Yafet (1963); Esteve-Paredes and Palacios (2023)? We recall that the discussion of Bloch electrons’ position is entirely based on an equilibrium picture, and the equilibrium location of an electron in the lattice is indeed undefined. Nevertheless, out of equilibrium an electric field in general displaces an electron’s center of mass away from its equilibrium location, generating a dipole. Focusing on the intrinsic case, in which electric-field induced changes to the density matrix are purely band off-diagonal, this dipole represents a DC version of the inter-band polarization known in non-linear optics Sipe and Shkrebtii (2000); King-Smith and Vanderbilt (1993). This suggests one mechanism behind the OHE as the generation of an inter-band polarization by the applied electric field: the electric field generates a dipole which then rotates about the centre of mass, generating an OAM, which is convected along with the electron. This is the physics behind $\Delta j_{1}$. In incorporating the band-diagonal position matrix elements one must ensure that no part of the final result depends on the equilibrium centre of mass of the electron. Our boundary conditions take this into account.
Building on this insight, if the OAM is regarded as an operator, one must all for this operator to have inter-band matrix elements, and these can contribute to OAM transport, in the same way as the inter-band position and velocity matrix elements are responsible for the equilibrium OAM. Such band off-diagonal matrix elements of the OAM operator reflect inter-band coherence induced by an electric field. It is these matrix elements that are primarily responsible for $\Delta j_{2}$. Next, once the position and velocity are treated as full operators, one must also account for the fact that they do not commute, and this leads directly to $\Delta j_{3}$.
We have focussed on the Bloch picture of electrons in crystals. Different insights into the OAM emerge from the Wannier and Bloch formulations of the modern theory Rhonald Burgos Atencia and Culcer (2024); Cysne et al. (2022). In the Wannier picture, the OAM involves intra-atomic and inter-atomic contributions Pezo et al. (2022, 2023); Busch et al. (2023); Xuan and Quek (2020), yet an exact correspondence between individual terms in the OAM between the two approaches is not always possible Rhonald Burgos Atencia and Culcer (2024). Nevertheless, since an exact correspondence exists between the Bloch and Wannier bases, the fundamental considerations outlined in this work will be applicable to Wannier studies as well.
@@
We stress that our work is concerned with the evaluation of the current, not its definition. In the long run the same fundamental issues will need to be considered for the OHE as for the spin-Hall effect. One is that the orbital Hall current, as defined, can be nonzero in equilibrium in the same way as the spin current Rashba (2003). In fact all the four contributions to $j_{L}$ could be separately nonzero in equilibrium, although they do vanish for all the effective models we have investigated – CuMnAs, bulk topological insulators, and massive Dirac fermions. Secondly, the OAM itself may not be conserved in an electric field, as we recently showed Atencia et al. (2024), and one may eventually need to consider a proper OH current in analogy with the proper spin current. Thirdly, the only observable is the OAM, and the relationship between the current and the OAM accumulation at the boundary will have to be determined. Nevertheless, no matter which approach is taken, the fundamental problem of the evaluation of $j_{L}$ is relevant. In this context, to examine OAM conservation in our chosen example we calculated the torque term $\langle\partial_{t}\hat{L}_{\alpha}\rangle$ for the CuMnAs tight-binding model Eq. (6) following the prescription of Ref. Atencia et al. (2024), and found $\langle\partial_{t}\hat{L}_{\alpha}\rangle=0$. This implies the OAM is conserved in this model even in an electric field.
Figure 3: The intrinsic $j_{L}$, extrinsic $j_{\rm dis}$, and total OHE for CuMnAs, with the same parameters as in Fig. 1.
Disorder. Non-equilibrium effects can be classified as intrinsic and extrinsic depending on whether they originate in the band structure or disorder. Terms to zeroth order in the disorder strength play crucial roles in Hall effects Khaetskii (2006); Gradhand et al. (2010); Ferreira et al. (2014); Cullen and Culcer (2023); Liu and Culcer (2024). In the density matrix formalism such corrections are incorporated into an anomalous driving term Culcer et al. (2017) which results in a correction to $\rho^{mn}_{E{\bm{k}}}$ in Eq. 4. In a recent work we have argued that this disorder correction dominates the OHE in massless Dirac fermion systems Liu and Culcer (2024), where our calculation was based on the conventional evaluation of the OHE. Here we estimate the extrinsic contribution to the full orbital current in the relaxation-time approximation, where the inverse relaxation time serves as a measure of the disorder strength. The extrinsic OHE is zeroth-order in the relaxation time, that is, the same order as the intrinsic OHE Liu and Culcer (2024). The relaxation time approximation does not take into account the electrical field corrected scattering integral Atencia et al. (2022), which is extremely laborious. We note also that the Eq. (6) is a 4-band model, which is considerably more complex than the two-band tilt Dirac model of Ref. Liu and Culcer (2024). Our results are shown in Fig. 3. We find that the extrinsic contribution (i) is much larger than the conventional OHE, in agreement with Ref. Liu and Culcer (2024), and (ii) is of the same order of magnitude as intrinsic contribution to the total $j_{L}$, including the quantum correction.
@@
Conclusion. We have presented a full quantum mechanical evaluation of the orbital current and orbital Hall effect, demonstrating that the existence of a quantum correction that is missed in conventional evaluations. This correction stems from the band-diagonal terms in the position and velocity operators, as well as the non-commutativity of the position and velocity operators. In the tight-binding Hamiltonian of tetragonal CuMnAs the quantum correction is by far the dominant contribution to the OHE and changes its sign.
@@
Acknowledgments. This work is supported by the Australian Research Council Centre of Excellence in Future Low-Energy Electronics Technologies, project number CE170100039. We acknowledge enlightening discussions with Naoto Nagaosa, Guang-Yu Guo, Hsin Lin, Mikhail Titov, and Tatiana Rappoport.

### v2

@@ Abstract
What is the physical meaning of the quantum corrections to the OHE? We recall that the discussion of Bloch electrons’ position is entirely based on an equilibrium picture, and the equilibrium location of an electron in the lattice is undefined. Nevertheless, out of equilibrium an electric field in general displaces an electron’s center of mass away from its equilibrium location, generating a dipole. In the intrinsic case, electric-field induced changes to the density matrix are purely band off-diagonal, this dipole represents a DC version of the inter-band polarization known in non-linear optics [93, 94]. This suggests one mechanism behind the OHE as the generation of an inter-band polarization by the applied electric field: the electric field generates a dipole which then rotates about the centre of mass, generating an OAM, which is convected along with the electron. This is the physics behind $\Delta j_{1}$  [^2] In incorporating the band-diagonal position matrix elements one must ensure that no part of the final result depends on the equilibrium centre of mass of the electron. Our boundary conditions take this into account.. Next, the OAM operator itself has inter-band matrix elements. These do not contribute to the equilibrium OAM expectation value, because the equilibrium density matrix is diagonal in the band index. Nevertheless, the inter-band matrix elements of the OAM operator do contribute to OAM transport – they contribute to the OHE. It is these matrix elements that are primarily responsible for $\Delta j_{2}$. Finally, once the position and velocity are treated as full operators, one must also account for the fact that they do not commute, and this leads directly to $\Delta j_{3}$. In equilibrium, this commutator can be related to the effective mass,[96] yet out of equilibrium the resulting expression is considerably more complex. Disentangling the different contributions experimentally is challenging. However, in the case of CuMnAs the predictions differ by three orders of magnitude, so a comparison with experimental results, once they become available, can be made unambiguously.
@@
We stress that our work is concerned with the evaluation of the current, not its definition. In the long run the same fundamental issues will need to be considered for the OHE as for the spin-Hall effect. One is that the orbital Hall current, as defined, can be nonzero in equilibrium in the same way as the spin current [97]. In fact all the four contributions to $j_{L}$ could be separately nonzero in equilibrium, although they do vanish for all the effective models we have investigated – CuMnAs, bulk topological insulators, and massive Dirac fermions. Secondly, the OAM itself may not be conserved in an electric field, as we recently showed [98], and one may eventually need to consider a proper OH current in analogy with the proper spin current. Thirdly, the only observable is the OAM, and the relationship between the current and the OAM accumulation at the boundary will have to be determined. Nevertheless, whichever approach is taken, the fundamental problem of the evaluation of $j_{L}$ is relevant.
@@
Conclusion. We have presented a full quantum mechanical evaluation of the OHE, demonstrating that the existence of quantum corrections that have been missed in conventional evaluations. These corrections stem from the band-diagonal terms in the position and velocity operators, as well as the non-commutativity of the position and velocity operators. In the models considered the quantum corrections are at least as large as the conventional terms in the OHE, and in CuMnAs they overwhelm the conventional result.
@@
Acknowledgments – This work is supported by the Australian Research Council Centre of Excellence in Future Low-Energy Electronics Technologies, project number CE170100039. JHC acknowledges support from an Australian Government Research Training Program (RTP) Scholarship. We acknowledge enlightening discussions with Rhonald Burgos, Qian Niu, Naoto Nagaosa, Guang-Yu Guo, Hsin Lin, Mikhail Titov, Aires Ferreira, Tatiana Rappoport, and Giovanni Vignale.